12/7y2 * y5 * 14y - (4/5)y2 * 5y2

1
Expert's answer
2021-05-17T04:29:48-0400

"\Rightarrow\dfrac{12}{7y^2}\times y^5\times14y-\dfrac{4}{5}y^2\times5y^2"


"\Rightarrow 24 \dfrac{y^6}{y^2}-4y^4"


"\Rightarrow24 y^4-4y^4=20y^4"